A cable 20 feet long connects the top of a flagpole to a point on the ground that is 16 feet from the base of the pole. How tall is the flagpole? (A). 8ft. (B). 10ft. (C). 12ft. (C). 26ft
Answer by mananth(16949)

The rope,ground line and the pole form a right triangle
Leg 1 = 16 ft
Hypotenuse = 20 feet
Leg 2 is the distance from foot of pole to the point the rope is tied = x
20^2-16^2= x^2
400-256=x^2
144 = x^2
take the square root
x= 12 feet